A satellite is projected near the surface of Earth so that it has a certain speed that can orbit the Earth. Calculate the speed of the satellite and the time for 1 complete rotation around the Earth. (Given, ME = 6 x 1024 kg, RE = 6.38 x 106 m, G = 6.67 x 10-11 Nm2kg-2)

1
Expert's answer
2022-04-18T08:50:07-0400

(i) the speed of the satellite

"v=\sqrt{G\frac{M_E}{r_E}}\\
=\sqrt{6.67*10^{-11}*\frac{6*10^{24}}{6.38*10^6}}=7920\;\rm m/s"

(ii) the period of rotation

"T=\frac{2\pi r_E}{v}\\
=\frac{6.28*6.38*10^6}{7920}=5059\:\rm s=84\: min"